OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4).

OpenStudy (diyadiya):

Just plug in -4 in place of x

OpenStudy (diyadiya):

-4(-4)+5

OpenStudy (anonymous):

-4(-4)+5 16+5=21

OpenStudy (anonymous):

g(-4)=16+5=21

OpenStudy (anonymous):

So I got g(x)=-4x+5 g(-4)= -4(-4)+5

OpenStudy (diyadiya):

Yes right

OpenStudy (anonymous):

g(-4)=16+5 g(-4=21

OpenStudy (diyadiya):

Right! :)

OpenStudy (anonymous):

did I solve right?

OpenStudy (diyadiya):

Absolutely right!